OpenStudy (anonymous):

Find the number of solutions for the given equation for 0° ≤ θ ≤ 360°. sin2x - 2sinx - 3 = 0

OpenStudy (campbell_st):

is the equation\[\sin^2 x - 2sinx - 3=0\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Its a quadratic in "sin(x)". Solve for the roots of the quadratic. Then, use inverse of sine to find x.

OpenStudy (campbell_st):

you need to factorise the quadratic and solve.... but there is only 1answer to the problem x = 270

OpenStudy (anonymous):

This is supposed to be a Trigonometric Equation.

OpenStudy (anonymous):

It is a qudratic equation with trigonometric function in its terms.

OpenStudy (campbell_st):

the factorisation is (sin x -3)(sinx + 1) = 0

OpenStudy (anonymous):

Since sin(x) can't be 3 ever, it only has one solution, which is sin(x) = -1 => x = 270 degrees.